# Linear cost and exponentially convergent approximation of Gaussian Matérn processes

### v1

@@ 4.4 Process approximation
As can be seen in the previous section, no alternative method is close to the proposed method in terms of accuracy of the covariance approximation. The only methods that are close to the proposed method in terms of prediction accuracy are the nnGP and FEM methods. One important difference of the nnGP method is that the approximation is strongly dependent on the support points, which typically is chosen as the observation locations in applications (Datta et al., 2016). In this section, we illustrate that this results in a poor approximation of the stochastic process, even though the approximation of the finite dimensional distribution at the support points is accurate. To this end, we consider the task of computing a joint confidence band for the latent process. Specifically, we assume that we have data ${\boldsymbol{\mathrm{y}}}=[y_{1},y_{2},\dotsc,y_{n}]$ generated as $y_{i}|u(\cdot)\sim\mathcal{N}(u(t_{i}),\sigma_{e}^{2}).$ where $t_{i}$ are $250$ locations sampled uniformly at random in the interval $(0,10)$, and $u$ is a centered Gaussian process with the Matérn covariance (1). Based on these observations, we perform prediction to $n$ evenly spaced locations in the interval, and we use the excursions package (Bolin & Lindgren, 2018) to compute the upper and lower bounds, $c_{l}(s)$ and $c_{u}(s)$, of a joint prediction interval so that
$P(c_{l}(s)<u(s)<c_{u}(s)|{\boldsymbol{\mathrm{y}}},s\in\{p_{1},\ldots,p_{n}\})=0.9.$
@@
Then, for nnGP, FEM, and the proposed method, we compute the approximate coverage probabilities $\tilde{p}=\tilde{P}(c_{l}(s)<u(s)<c_{u}(s)|{\boldsymbol{\mathrm{y}}},s\in\{p_{1},\ldots,p_{n}\}),$ where $\tilde{P}(\cdot|\cdot)$ denote the posterior probability under the approximate model. If the process approximation is accurate, we expect that $\tilde{p}\approx 0.9$.
We choose $\nu=1$ and three values, $\rho=0.5,1$ and $2$, of the practical correlation range $\rho=\nicefrac{{\sqrt{8\nu}}}{{\kappa}}$. We perform this exercise for different values of $n$, while keeping the number of observations fixed at $250$. For each value of $n$, the order $m$ of the nnGP approximation is chosen so that the computational cost for evaluating the posterior mean at the prediction locations is the same as for the proposed method, and the support points for the nnGP approximation is kept fixed at the observation locations. For $n<1000$, we use the calibration cost from $n=1000$, as the observed cost in smaller cases primarily reflects computational overheads rather than the intrinsic complexity of the methods.
The error $|\tilde{p}-0.9|$ for the methods are shown as functions of $n$ in Figure 3. As expected, the error of the nnGP method increases quickly with $n$, whereas it remains more stable for the proposed approximation. This supports the claim that the proposed approximation is a better process approximation than the nnGP method. Further, the approximation is also better than the corresponding (calibrated) ones from the FEM method.
@@
Figure 3: Absolute posterior probability errors on the interval $[0,10]$, based on $250$ observations and $N$ additional locations for $\nu=1$, $m$ from $3$ to $6$, and three different practical correlation ranges $\rho$.
@@ 5 Discussion
Our experimental results demonstrate that the proposed method outperforms state-of-the-art alternatives, offering significant improvements in accuracy. The method has been implemented in the rSPDE package, which is compatible with popular tools such as R-INLA (Lindgren & Rue, 2015) and inlabru (Bachl et al., 2019), facilitating seamless integration into Bayesian hierarchical models. Further, all codes and a Shiny application can be found at https://github.com/vpnsctl/MarkovApproxMatern/.
The Shiny application also contains further comparisons for other choices of the number of prediction locations $n$, the number of observation locations $n_{obs}$, and for different practical correlation ranges of the true process. However, the general conclusions for these different choices align closely with those presented above. The main difference is that, for prediction tasks with a very small practical correlation range and an equal number of prediction and observation locations, nnGP tends to outperform all other methods, including the rational approximation. This scenario is particularly favorable for nnGP, as all locations coincide with support points. Additionally, when the number of prediction locations exceeds the number of observation locations, the rational approximation outperforms all methods, except for $\nu>1.5$, where nnGP achieves the best performance in prediction tasks.
The two most competitive methods besides the proposed method were nnGP and FEM. The nnGP model provides accurate predictions, particularly for lower values of the smoothness parameter $\nu$, as illustrated in Figure 2. However, it falls short in terms of process approximation. Specifically, the covariance errors associated with nnGP are significantly larger than those from the rational approximation method, as shown in Figure 1. Additionally, the probability approximations deteriorate markedly when observations are not contained in the support points, evidenced by Figure 3. This limitation suggests that while nnGP may provide good results in contexts focused on prediction, such as those commonly found in machine learning, it is less competitive in broader statistical applications which require accurate process approximations. The main advantage of the proposed method compared to the FEM method is a generally higher accuracy and stability, and the fact that no mesh needs to be chosen.
@@
$L^{-2\beta}\approx L^{-\lfloor 2\beta\rfloor}\left(\sum_{i=1}^{m}r_{i}(L-p_{i}I)^{-1}+kI\right),$ (13)
where $\{r_{i}\}_{i=1}^{m}$, $\{p_{i}\}_{i=1}^{m}$, and $k$ are real numbers, and $I:L_{2}(\mathcal{D})\to L_{2}(\mathcal{D})$ is the identity operator. If $p_{i}\leq 0$, $r_{i}>0$, and $k>0$ for all $i=1,\dots,m$, and suitable conditions on $L$ hold, this approximation expresses $L^{-2\beta}$ as a linear combination of covariance operators. We will refer to this approach as the covariance-based approach.
By computing the finite-dimensional distributions corresponding to the operators $L^{-\lfloor 2\beta\rfloor}$ and $L^{-\lfloor 2\beta\rfloor}(L-p_{i}I)^{-1}$ for $i=1,\dots,m$, we can approximate the finite-dimensional distributions of $L^{-2\beta}$ and thus the solution to (12). Assume further that these covariance operators are induced by covariance functions. Let $\varrho_{\beta}(\cdot,\cdot)$ denote the covariance function associated with $L^{-2\beta}$, $\widehat{\varrho}_{\beta}(\cdot,\cdot)$ the covariance function corresponding to $kL^{-\lfloor 2\beta\rfloor}$, and $\widehat{\varrho}_{\beta,i}(\cdot,\cdot)$ the covariance function associated with $r_{i}L^{-\lfloor 2\beta\rfloor}(L-p_{i}I)^{-1}$ for $i=1,\dots,m$. Then, (13) implies that
@@
which corresponds exactly to the expression in Proposition 2.1, demonstrating the equivalence between the spectral and covariance-based approaches.
@@
Appendix A LDL Algorithm and proof of Proposition 3.4
In this section, we explain how to construct the matrices ${\boldsymbol{\mathrm{L}}}_{i}$ and ${\boldsymbol{\mathrm{D}}}_{i}$ of Proposition 3.4. To simplify the notation, let $p$ denote the size of the vector ${\boldsymbol{\mathrm{u}}}_{i,j}$ (which is either $\max(\lfloor\alpha\rfloor,1)$ if $i=0$ or $\lceil\alpha\rceil$ if $i>0$). Introduce the matrices ${\boldsymbol{\mathrm{B}}}^{1}={\boldsymbol{\mathrm{L}}}_{1,1}$ and ${\boldsymbol{\mathrm{B}}}^{k}=[{\boldsymbol{\mathrm{L}}}_{k,k-1}\,{\boldsymbol{\mathrm{L}}}_{k,k}],k>1,$ corresponding to the part of ${\boldsymbol{\mathrm{L}}}_{i}$ that is related to the $k$th location. Similarly, let ${\boldsymbol{\mathrm{F}}}^{k}$ denote the diagonal matrix corresponding to the part of ${\boldsymbol{\mathrm{D}}}^{-1}$ that is related to the $k$th location. That is, ${\boldsymbol{\mathrm{D}}}=\text{diag}({\boldsymbol{\mathrm{F}}}^{1},\ldots,{\boldsymbol{\mathrm{F}}}^{n})^{-1}.$ These are low-dimensional matrices with a dimension $2p\times p$ and $p\times p$, respectively. Further, introduce the covariance matrices ${\boldsymbol{\mathrm{\Sigma}}}^{1}={\boldsymbol{\mathrm{r}}}_{i}(t_{1},t_{1})$ and
@@ Appendix B Collected proofs
We start with a simple technical lemma needed for the proof of Theorem 2.2.
Lemma B.1.
@@
Proof B.2.
We have, by the change of variables $u=x-y$, that
@@
Our idea is to use the exponential convergence of the best rational approximation with respect to the $L_{\infty}$ norm. To such an end, Stahl (2003, Theorem 1), gives us that for every $\alpha\in(0,1)$, there exist polynomials of degree $m$, $p_{m}(\cdot)$ and $q_{m}(\cdot)$ such that the following exponential bound holds:
$\sup_{x\in[0,1]}\left|x^{\alpha}-\frac{p_{m}(x)}{q_{m}(x)}\right|\leq Ce^{-2\pi\sqrt{\alpha m}},$ (15)
@@
where $C>0$ is a constant that only depends on $\alpha$. Now, let $g(w)=(1+\kappa^{-2}w^{2})^{-1}$ and observe that for every $w\in\mathbb{R}$, we have $g(w)\in[0,1]$. Further, let $\alpha=\nu+0.5$ and $f^{\alpha}(\cdot)$ be the spectral density of a stationary Gaussian process with a Matérn covariance function (1). Then, $f^{\alpha}=\kappa^{-2\alpha}(g(w))^{\alpha}$ and for any $\alpha>1/2$ we have that
$\sup_{w\in\mathbb{R}}\left|f^{\alpha}(w)-f_{m}^{\alpha}(w)\right|$ $=\kappa^{-2\alpha}\sup_{w\in\mathbb{R}}\left|(g(w))^{\alpha}-g(w)^{\lfloor\alpha\rfloor}\frac{p_{m}(g(w))}{q_{m}(g(w))}\right|=\kappa^{-2\alpha}\sup_{x\in[0,1]}\left|x^{\alpha}-x^{\lfloor\alpha\rfloor}\frac{p_{m}(x)}{q_{m}(x)}\right|$
$=\kappa^{-2\alpha}\sup_{x\in[0,1]}x^{\lfloor\alpha\rfloor}\left|x^{\{\alpha\}}-\frac{p_{m}(x)}{q_{m}(x)}\right|\leq Ce^{-2\pi\sqrt{\alpha m}},$
@@
where we used (15) and the fact that $0\leq x\leq 1$. Now, let $I=[a,b]\subset\mathbb{R}$ be an interval in $\mathbb{R}$ and $1_{I}:\mathbb{R}\rightarrow\mathbb{R}$ denote the indicator function such that $11_{I}(x)=$ 1 if $x\in I$ and $01_{I}(x)=$ 0 otherwise. We then have by Plancherel’s theorem (e.g. McLean, 2000, Corollary 3.13), the convolution theorem (e.g. McLean, 2000, p.73), Lemma B.1, and the bound above that
$\|r_{m}^{\alpha}-r^{\alpha}\|_{L_{2}(I\times I)}^{2}$ $\leq(b-a)\|\varrho_{m}^{\alpha}-\varrho^{\alpha}\|_{L_{2}(a-b,b-a)}^{2}=(b-a)\|(\varrho_{m}^{\alpha}-\varrho^{\alpha})1_{[a-b,b-a)]}\|_{L_{2}(\mathbb{R})}^{2}$
@@
$=(b-a)\left\|(f_{m}^{\alpha}-f^{\alpha})\ast 2(b-a)\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
$\leq(b-a)\sup_{w\in\mathbb{R}}\left(f^{\{\alpha\}}(w)-\frac{p_{m}(w)}{q_{m}(w)}\right)^{2}\left\|2(b-a)\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
$\leq(b-a)\widehat{C}e^{-2\pi\sqrt{\{\alpha\}m}},$
where the constant $\widehat{C}$ only depends on $\alpha,\kappa$ and $\sigma$, and $\ast$ denotes the convolution. This proves (5).
@@
$\|\varrho^{\alpha}-\varrho_{m}^{\alpha}\|_{H^{s}(\mathbb{R})}$ $=\|(1+w^{2})^{\nicefrac{{s}}{{2}}}(f^{\alpha}-f_{m}^{\alpha})\|_{L_{2}(\mathbb{R})}\leq\sup_{w\in\mathbb{R}}\left|f^{\{\alpha\}}(w)-\frac{p_{m}(w)}{q_{m}(w)}\right|\|f^{\lfloor\alpha\rfloor}(1+w^{2})^{\nicefrac{{s}}{{2}}}\|_{L_{2}(\mathbb{R})}$
$\leq\widetilde{C}e^{-2\pi\sqrt{\{\alpha\}m}}\|(1+w^{2})^{s/2-\lfloor\alpha\rfloor}\|_{L_{2}(\mathbb{R})},$

### v2

@@ 4.5 Process approximation
As observed in Section 4.3, no alternative method comes close to the proposed method in terms of the stability and accuracy of the covariance approximation. However, in terms of prediction accuracy, the nnGP and FEM methods exhibit similar performance in certain scenarios. Specifically, for the FEM method, similar accuracies are observed for small values of $\nu$, such as $\nu<1$ for order 3 and $\nu<0.5$ for order 5. Also, for certain cases, the nnGP method outperformed the proposed method in terms of prediction accuracy. However, one important difference of the nnGP method is that the approximation is strongly dependent on the support points, which typically is chosen as the observation locations in applications (Datta et al., 2016). In this section, we illustrate that this results in a poor approximation of the stochastic process, even though the approximation of the finite dimensional distribution at the support points is accurate. To this end, we consider the task of computing a joint confidence band for the latent process in two different scenarios. In both scenarios, we assume that we have data $\mathbf{y}=[y_{1},y_{2},\dotsc,y_{n}]$ generated as $y_{i}|u(\cdot)\sim\mathcal{N}(u(t_{i}),\sigma_{e}^{2}),$ where $u$ is a centered Gaussian process with the Matérn covariance (1). As before, we consider the two noise levels $\sigma_{e}=0.1$ and $\sigma_{e}=\sqrt{0.1}$. Based on these observations, we perform prediction at locations $p_{1},\ldots,p_{n}$, and use the excursions package (Bolin and Lindgren, 2018) to compute the upper and lower bounds, $c_{l}(s)$ and $c_{u}(s)$, of a joint confidence band so that $P(c_{l}(s)<u(s)<c_{u}(s)|\mathbf{y},s\in\{p_{1},\ldots,p_{n}\})=0.9.$
In Scenario 1, we consider a forecasting setting where we construct a regular mesh of $1501$ points in the interval $[0,15]$. The first $1001$ points, evenly spaced in $[0,10]$, serve as observation locations. For prediction, we use all observation locations plus the next ${n\in\{0,1,\ldots,10,20,30,40,50,75,100,125,\ldots,500\}}$ consecutive points from the mesh as prediction locations. In Scenario 2, we instead use $250$ observation locations randomly sampled uniformly in the interval $[0,10]$, with a minimum spacing of $10^{-3}$ between locations ensured through resampling if needed, to ensure stability for nnGP. For for an increasing sequence of values $n$ between 00 and $3000$, we consider $n$ evenly spaced prediction locations in the interval. The combined set of observation and prediction locations is processed to maintain the minimum spacing requirement while preserving all observation locations. To account for the randomness in the selection of observation locations, we repeat this scenario $10$ times and average the results.
@@
For each scenario, we compute the approximate coverage probabilities for nnGP, FEM, and the proposed method: $\tilde{p}=\tilde{P}(c_{l}(s)<u(s)<c_{u}(s)|\mathbf{y},s\in\{p_{1},\ldots,p_{n}\}),$ where $\tilde{P}(\cdot|\cdot)$ denotes the posterior probability under the approximate model. If the process approximation is accurate, we expect that $\tilde{p}\approx 0.9$. Additionally, we evaluate the accuracy using the Kullback-Leibler (KL) divergence between the approximate and true posterior distributions.
@@
Figure 6: Posterior probability errors and KL divergence under Scenario 1 with $\sigma_{e}=0.1$ for $\nu=1$, orders $m$ from $3$ to $6$, and three different practical correlation ranges $\rho$.
We choose $\nu=1$ and three values of the practical correlation range: $\rho=0.5,1,$ and $2$. For each scenario and each value of $n$, the order $m$ of the nnGP approximation is chosen so that the computational cost for evaluating the posterior mean at the prediction locations is the same as for the proposed method, and the support points for the nnGP approximation is kept fixed at the observation locations. For $n<1000$, we use the calibration cost from $n=1000$, as the observed cost in smaller cases primarily reflects computational overhead.
The error $0.9-\tilde{p}$ and KL divergences for the rational, nnGP and FEM methods are shown as functions of $n$ for Scenario 1 and noise level $\sigma_{e}=0.1$ in Figure 6 and for Scenario 2 and noise level $\sigma_{e}=\sqrt{0.1}$ in Figure 7. The results for Scenario 1 with $\sigma_{e}=\sqrt{0.1}$ and Scenario 2 with $\sigma_{e}=0.1$ are very similar, so we do not include them here. However, the results for those cases can be seen in the accompanying Shiny app.
As expected, the error of the nnGP method is small for $n=0$ but increases quickly with $n$ and has poor performance both in terms of probability approximation and KL divergence even for a very small number of prediction locations. The results for the proposed approximation are much more stable, which supports the claim that the proposed approximation is a better process approximation than the nnGP method. Further, the approximation is also better than the corresponding (calibrated) ones from the FEM method.
Figure 7: Posterior probability errors and KL divergence under Scenario 2 with $\sigma_{e}=\sqrt{0.1}$ for $\nu=1$, orders $m$ from $3$ to $6$, and three different practical correlation ranges $\rho$.
As a final study, we consider Scenario 1, in which we do forecasting, and compute the KL divergence between the true distribution of the field $(u(p_{1}),\ldots,u(p_{n}))$, where $u(\cdot)$ is a Gaussian process with the Matérn covariance function 1, and their approximations by the rational, nnGP and FEM methods. We follow the approach of Scenario 1 and consider the first $1001$ points as support points for nnGP. We then compute the KL divergence between the true distribution of the field and the approximations by the different methods. The results are shown in Figure 8, and we observe a similar situation as in the posterior distributions study. The proposed method outperforms the FEM method consistently, and the nnGP method has good performance if all locations are support points. However, the nnGP error increases rapidly as the number of non-support points increases, and as soon as we have a few non-support points, the proposed method also outperforms nnGP.
Remark 12
We did not include Scenario 2 in the prior distribution study because the nnGP method is not stable for this scenario. The reason for this instability is that the nnGP method is highly dependent on the ordering of the support points and prediction points as shown in (Schäfer et al., 2021, Section 4.2.1). For the posterior calculations we were able to overcome this instability by using a Markov property for performing prediction for nnGP (Datta et al., 2016), which stabilized the posterior distributions. However, as seen in Figures 6 through 8, it is unlikely that the conclusions would be different for prior distributions under Scenario 2.
Figure 8: KL divergences (in log scale) under Scenario 1 for $N$ non-support locations, $\nu=1$, orders $m$ from $1$ to $6$, and three different practical correlation ranges $\rho$.
@@ 5 Discussion
Our experimental results demonstrate that the proposed method outperforms state-of-the-art alternatives, offering significant improvements in accuracy. The method has been implemented in the rSPDE package, which is compatible with popular tools such as R-INLA (Lindgren and Rue, 2015) and inlabru (Bachl et al., 2019), facilitating seamless integration into Bayesian hierarchical models.
The two most competitive methods besides the proposed method were nnGP and FEM. The nnGP model provides accurate predictions, particularly for lower values of the smoothness parameter $\nu$, as illustrated in Figure 4. However, it falls short in terms of process approximation. Specifically, the covariance errors associated with nnGP are significantly larger than those from the rational approximation method, as shown in Figure 3. Additionally, the probability approximations deteriorate markedly when observations are not contained in the support points, evidenced by Figures 6 through 8. This limitation suggests that while nnGP may provide good results in contexts focused on prediction, such as those commonly found in machine learning, it is less competitive in broader statistical applications which require accurate process approximations. The main advantage of the proposed method compared to the FEM method is a generally higher accuracy and stability (in particular for large values of $\nu$), and the fact that no mesh needs to be chosen.
To ensure a fair comparison, we assumed all parameters to be known throughout the simulation studies, providing a clear baseline using the true covariance function. We expect similar results also in cases with unknown parameters, such as when performing likelihood-based statistical inference, as iterative numerical procedures rely on intermediate approximations of the true covariance. However, this remains to be investigated.
@@
$L^{-2\beta}\approx L^{-\lfloor 2\beta\rfloor}\left(\sum_{i=1}^{m}c_{i}(L-p_{i}I)^{-1}+kI\right),$ (16)
where $\{c_{i}\}_{i=1}^{m}$, $\{p_{i}\}_{i=1}^{m}$ and $k$ are real numbers, satisfying $p_{i}\leq 0$, $c_{i}>0$, and $k>0$ by Proposition 3, and $I:L_{2}(\mathcal{D})\to L_{2}(\mathcal{D})$ is the identity operator. If suitable conditions on $L$ hold, this approximation expresses $L^{-2\beta}$ as a linear combination of covariance operators. We will refer to this approach as the covariance-based approach.
By computing the finite-dimensional distributions corresponding to the operators $L^{-\lfloor 2\beta\rfloor}$ and $L^{-\lfloor 2\beta\rfloor}(L-p_{i}I)^{-1}$ for $i=1,\dots,m$, we can approximate the finite-dimensional distributions of $L^{-2\beta}$ and thus the solution to (15). Assume further that these covariance operators are induced by covariance functions. Let $\varrho_{\beta}(\cdot,\cdot)$ denote the covariance function associated with $L^{-2\beta}$, $\widehat{\varrho}_{\beta}(\cdot,\cdot)$ the covariance function corresponding to $kL^{-\lfloor 2\beta\rfloor}$, and $\widehat{\varrho}_{\beta,i}(\cdot,\cdot)$ the covariance function associated with $c_{i}L^{-\lfloor 2\beta\rfloor}(L-p_{i}I)^{-1}$ for $i=1,\dots,m$. Then, (16) implies that
@@
which corresponds exactly to the expression in Proposition 5, demonstrating the equivalence between the spectral and covariance-based approaches.
@@
Another advantage with the covariance-based formulation is that is that it can be combined with FEM approximations, and because by taking advantage of the fact that (17) is stationary, one could use this to avoid having to rely on boundary extensions to remove boundary effects for the FEM approach. As this is a common problem with the FEM approach, which increases the computational cost, this is an interesting topic to investigate in future work.
Appendix A LDL Algorithm and proof of Proposition 9
In this section, we explain how to construct the matrices ${\boldsymbol{\mathrm{L}}}_{i}$ and ${\boldsymbol{\mathrm{D}}}_{i}$ of Proposition 9. To simplify the notation, let $p$ denote the size of the vector ${\boldsymbol{\mathrm{u}}}_{i,j}$ (which is either $\max(\lfloor\alpha\rfloor,1)$ if $i=0$ or $\lceil\alpha\rceil$ if $i>0$). Introduce the matrices ${\boldsymbol{\mathrm{B}}}^{1}={\boldsymbol{\mathrm{L}}}_{1,1}$ and ${\boldsymbol{\mathrm{B}}}^{k}=[{\boldsymbol{\mathrm{L}}}_{k,k-1}\,{\boldsymbol{\mathrm{L}}}_{k,k}],k>1,$ corresponding to the part of ${\boldsymbol{\mathrm{L}}}_{i}$ that is related to the $k$th location. Similarly, let ${\boldsymbol{\mathrm{F}}}^{k}$ denote the diagonal matrix corresponding to the part of ${\boldsymbol{\mathrm{D}}}^{-1}$ that is related to the $k$th location. That is, ${\boldsymbol{\mathrm{D}}}=\text{diag}({\boldsymbol{\mathrm{F}}}^{1},\ldots,{\boldsymbol{\mathrm{F}}}^{n})^{-1}.$ These are low-dimensional matrices with a dimension $2p\times p$ and $p\times p$, respectively. Further, introduce the covariance matrices ${\boldsymbol{\mathrm{\Sigma}}}^{1}={\boldsymbol{\mathrm{r}}}_{i}(t_{1},t_{1})$ and
@@ Appendix B Collected proofs
We start with a simple technical lemma needed for the proof of Theorem 1.
Lemma 13
@@
Proof  We have, by the change of variables $u=x-y$, that
[…]
@@
Proof [of Theorem 1] Our idea is to use the exponential convergence of the best rational approximation with respect to the $L_{\infty}$ norm. To such an end, Stahl (2003, Theorem 1), gives us that for every $\alpha\in(0,1)$, there exist polynomials of degree $m$, $p_{m}(\cdot)$ and $q_{m}(\cdot)$ such that the following exponential bound holds:
$\sup_{x\in[0,1]}\left|x^{\alpha}-\frac{p_{m}(x)}{q_{m}(x)}\right|\leq C_{\alpha}e^{-2\pi\sqrt{\alpha m}},$ (18)
@@
where $C_{\alpha}>0$ is a constant that only depends on $\alpha$. Now, let $g(w)=(1+\kappa^{-2}w^{2})^{-1}$ and observe that for every $w\in\mathbb{R}$, we have $g(w)\in[0,1]$. Further, let $\alpha=\nu+0.5$ and $f_{\alpha}(\cdot)$ be the spectral density of a stationary Gaussian process with a Matérn covariance function (1). Then, $f_{\alpha}(w)=A\sigma^{2}\kappa^{-2\alpha}(g(w))^{\alpha}$, where $A=\sqrt{2}\kappa^{2\nu}\Gamma(\nu+\nicefrac{{1}}{{2}})\Gamma(\nu)^{-1}$. Observe that we have the following decompositions:
$f_{\alpha}(w)=f_{\lfloor\alpha\rfloor}(w)(g(w))^{\{\alpha\}}\quad\hbox{and}\quad f_{m,\alpha}(w)=f_{\lfloor\alpha\rfloor}(w)\frac{p_{m}(g(w))}{q_{m}(g(w))}.$
We also have, for any $\alpha>1/2$, that
$\sup_{w\in\mathbb{R}}\left|(g(w))^{\{\alpha\}}-\frac{p_{m}(g(w))}{q_{m}(g(w))}\right|=\sup_{x\in[0,1]}\left|x^{\{\alpha\}}-\frac{p_{m}(x)}{q_{m}(x)}\right|\leq C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}},$
where we used (18) and that $0\leq x\leq 1$. Because $\sup_{w\in\mathbb{R}}|f_{\lfloor\alpha\rfloor}(w)|=|f_{\lfloor\alpha\rfloor}(0)|=A\sigma^{2}\kappa^{-2\alpha}$,
$\sup_{w\in\mathbb{R}}\left|f_{\alpha}(w)-f_{m,\alpha}(w)\right|$ $\leq\sup_{w\in\mathbb{R}}|f_{\lfloor\alpha\rfloor}(w)|\sup_{w\in\mathbb{R}}\left|(g(w))^{\{\alpha\}}-\frac{p_{m}(g(w))}{q_{m}(g(w))}\right|$
$\leq A\sigma^{2}\kappa^{-2\alpha}C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}}.$ (19)
@@
Now, let $I=[a,b]\subset\mathbb{R}$ be an interval in $\mathbb{R}$ and $1_{I}:\mathbb{R}\rightarrow\mathbb{R}$ denote the indicator function such that $11_{I}(x)=$ 1 if $x\in I$ and $01_{I}(x)=$ 0 otherwise. We then have by Plancherel’s theorem (e.g. McLean, 2000, Corollary 3.13), the convolution theorem (e.g. McLean, 2000, p.73), Lemma 13, and the bound (19) above, that
$\|r_{m,\alpha}-r_{\alpha}\|_{L_{2}(I\times I)}^{2}$ $\leq(b-a)\|\varrho_{m,\alpha}-\varrho_{\alpha}\|_{L_{2}(a-b,b-a)}^{2}=(b-a)\|(\varrho_{m,\alpha}-\varrho_{\alpha})1_{[a-b,b-a)]}\|_{L_{2}(\mathbb{R})}^{2}$
$=(b-a)\left\|(f_{m,\alpha}-f_{\alpha})\ast 2(b-a)\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
$\leq(b-a)\sup_{w\in\mathbb{R}}\left((g(w))^{\{\alpha\}}-\frac{p_{m}(w)}{q_{m}(w)}\right)^{2}\left\|2(b-a)\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
@@
$\leq 4(b-a)^{3}A^{2}\sigma^{4}\kappa^{-4\alpha}C_{\{\alpha\}}^{2}e^{-4\pi\sqrt{\{\alpha\}m}}\left\|\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
$=\pi(b-a)^{2}A^{2}\sigma^{4}\kappa^{-4\alpha}C_{\{\alpha\}}^{2}e^{-4\pi\sqrt{\{\alpha\}m}},$
where $\ast$ denotes the convolution and we used the fact that $\left\|\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}=\pi/(b-a)$. Let us now consider another strategy that helps us obtain another bound for $\alpha>1$, that gets better as $\alpha$ increases. First, observe that by Young’s convolution inequality (Bogachev, 2007, Theorem 3.9.4), we have $\|f\ast g\|_{L_{2}(\mathbb{R})}\leq\|f\|_{L_{1}(\mathbb{R})}\|g\|_{L_{2}(\mathbb{R})},$ which yields,
$\left\|f_{\lfloor\alpha\rfloor}(w)\ast\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}$ $\leq\|f_{\lfloor\alpha\rfloor}(w)\|_{L_{1}(\mathbb{R})}\left\|\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}$
$=A\sigma^{2}\kappa^{-2\alpha}M_{\lfloor\alpha\rfloor,\kappa}\sqrt{\frac{\pi}{b-a}},$
where $M_{0,\kappa}=\infty$ and $M_{n,\kappa}=\|(g(w))^{n}\|_{L_{1}(\mathbb{R})}=\kappa\pi\Gamma(2n-1)/(4^{n-1}\Gamma(n)^{2}),n\geq 1$. Therefore, by proceeding as in the first case, and using the above inequality, we have
$\|r_{m,\alpha}-r_{\alpha}\|_{L_{2}(I\times I)}^{2}$ $\leq(b-a)\left\|(f_{m,\alpha}-f_{\alpha})\ast 2(b-a)\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
$\leq 4(b-a)^{3}\sup_{w\in\mathbb{R}}\left((g(w))^{\{\alpha\}}-\frac{p_{m}(w)}{q_{m}(w)}\right)^{2}\left\|f_{\lfloor\alpha\rfloor}(w)\ast\frac{\sin((b-a)w)}{(b-a)w}\right\|_{L_{2}(\mathbb{R})}^{2}$
$\leq 4\pi(b-a)^{2}A^{2}\sigma^{4}\kappa^{-4\alpha}M_{\lfloor\alpha\rfloor,\kappa}^{2}C_{\{\alpha\}}^{2}e^{-4\pi\sqrt{\{\alpha\}m}}.$
Finally, by combining both inequalities above, we obtain
$\|r_{m,\alpha}-r_{\alpha}\|_{L_{2}(I\times I)}\leq 2\sqrt{\pi}(b-a)A\sigma^{2}\kappa^{-2\alpha}\min\{1,M_{\lfloor\alpha\rfloor,\kappa}\}C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}}.$
This proves (6).
@@
$\|\varrho_{\alpha}-\varrho_{m,\alpha}\|_{H^{s}(\mathbb{R})}$ $=\|(1+w^{2})^{\nicefrac{{s}}{{2}}}(f_{\alpha}-f_{m,\alpha})\|_{L_{2}(\mathbb{R})}$
@@
$\leq\sup_{w\in\mathbb{R}}\left|(g(w))^{\{\alpha\}}-\frac{p_{m}(w)}{q_{m}(w)}\right|\|f_{\lfloor\alpha\rfloor}(1+w^{2})^{\nicefrac{{s}}{{2}}}\|_{L_{2}(\mathbb{R})}$
$\leq A\sigma^{2}\kappa^{-2\alpha}C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}}\|(1+w^{2})^{s/2-\lfloor\alpha\rfloor}\|_{L_{2}(\mathbb{R})},$